OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7

OpenStudy (anonymous):

Alright, so what's the question?

OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)).

OpenStudy (anonymous):

Ok first lets find f(g(x)) any idea how to do that?

OpenStudy (anonymous):

Or can you please give an attempt however little it might be :)